OpenStudy (anonymous):

Find the maximum and minimum values of f(x,y)=xy on the ellipse 6x2+y2=5. Please help!!

OpenStudy (dumbcow):

first solve ellipse equation for y \[y = \sqrt{5-6x^{2}}\] substitute into f(x,y) \[f(x) = x \sqrt{5-6x^{2}}\] set derivative equal to 0 to find min/max \[f'(x) = \sqrt{5-6x^{2}}-\frac{6x^{2}}{\sqrt{5-6x^{2}}} = \frac{5-12x^{2}}{\sqrt{5-6x^{2}}}\] f'(x) = 0 when numerator is 0 \[5-12x^{2} = 0\] \[\ x = \pm \sqrt{\frac{5}{12}} \approx \pm 0.6455\] \[\max \rightarrow f(.6455) \approx 1.02\] \[\min \rightarrow f(-.6455) \approx -1.02\]